a mom is 3 times old as his daughter. 5 years ago their sum of ages were 50. how old is the mother
Answer by macston(5194)

d=daughter's age; m=mom's age=3d
If five years ago the sum of their ages were 50 years, the sum is currently 60 years.
d+m=60 yrs Substitute for m
d+3d=60yrs
4d=60yrs Divide each side by 4.
d=15yrs
m=3d=3(15yrs)=45yrs ANSWER: The mother is 45 years old.
CHECK:
(d-5yrs)+(m-5yrs)=50yrs
(15yrs-5yrs)+(45yrs-5yrs)=50yrs
10yrs+40yrs=50yrs
50yrs=50yrs